Essay. According to the U. S. Census Bureau, the total 2008 U.S. population was 303,824,640. The chart below summarizes the 2008 population for five U.S. States.
State 2008 Population
Missouri 5,911,605
Pennsylvania 12,448,279
Tennessee 6,214,888
Utah 2,736,424
Washington 6,549,224


SOURCE: U. S. Census Bureau

What is the probability that a randomly selected U.S. resident did not live in Utah? Show step by step work. Round solution to the nearest thousandth.


Answer by htmentor(1343)

It seems that the only relevant information in this problem is the population of Utah, and the total U.S. population.
The probability, P, that the person lives in Utah = Population of Utah divided by the U.S. population:
P = 2,736,424/303,824,640 = 0.009
The probability that the person DID NOT live in Utah = 1 - P = 1 - 0.009 = 0.991